OpenStudy (anonymous):

How many atoms are in 15.3 cm3 of mercury (at room temperature)

OpenStudy (aaronq):

find mass.. density=mass/volume find moles.. moles=mass/molar mass multiply by avogadros number to find the number of atoms

OpenStudy (anonymous):

Density = mass/volume you know the volume 15.3 cm^3 which equals 15.3 mL find the density of mercury (on a chart) then find the mass once you have the mass use n=m/MW (moles equals mass over molecular weight) find the number of moles (n) and multiply that by avogadros number to find the amount of particles (in your case, atoms) present.

OpenStudy (anonymous):

It doesnt work. The density of Hg is 15.546 g/cm3, i.e. according to your calculation: 15.546 x 15.3 = 237.85 237.85 x 6.002*10 23 = 1.4 *10 26, and this comes up as wrong when I enter it :/ Do I make anything wrong here?

OpenStudy (aaronq):

you forgot to convert the grams 237.85 Hg to moles of Hg

OpenStudy (anonymous):

Oh yeh thats right, thanks!